what is the de-broglie wavelength of electron of energy 180ev

Grade:12

5 Answers

Akash Kumar Dutta
11 years ago

apply lambda=hc/E...where E=180eV

Chetan Mandayam Nayakar
11 years ago

p2/2m=E, p=sqrt(2mE),λ=h/(2π)

Vikas TU
10 years ago

For de-broglie wavelength 

lamda = underoot 150/V 

where V is the Voltage

Lamda = underroot150/180

          = underroot5/6 is the ansr.

Abhishekh kumar sharma
10 years ago

we have e=hv    ... 1

and v=c\λ

substitue v in 1 to get e=hc\λ

then lambda =hc\e

Sukumar Adepu
10 years ago

9 * 10 power -11